OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, −3) and has a slope of 1/2 .. y − 6 = 1/2(x + 3) y = 1/2x − 6 y + 3 = 1/2(x − 6) x − 2y = 12

OpenStudy (anonymous):

The point-slope form of the equation the line that passes through the point (a,b) and has the slope m is y-b = m(x-a) Therefore, the equation the line that passes through the point (6, -3) and has the slope 1/2 is y - (-3) = 1/2 (x - 6) y + 3 = 1/2 (x-6)

OpenStudy (anonymous):

\[y-y _{1}= m(x-x _{1})\] where \[(x _{1},y _{1}) = (6,-3)\] and m = 1/2 so which do you think fits?